OpenStudy (newwar):

Which ordered pair is a solution to this equation? (x+6)y=9(x+6)y=9 (3, 1) (4, 1) (12, 2) (2, 1)

OpenStudy (newwar):

@DanJS

OpenStudy (3mar):

can we finish the last one firstly?

OpenStudy (newwar):



OpenStudy (newwar):

ok

OpenStudy (3mar):

It is only single expression, Dan. He just made a copy mistake.

OpenStudy (danjs):

(x + 6) *y = 9 Test each point to see if that equation is true or not. (x,y) = (3,1), x=3, y=1 (3 + 6)* 1 = 9 9 = 9 true

OpenStudy (3mar):

\[(x+6)y=9\]